choose an article on nutrition topics from any reliable sources (journal/science daily/nutrition magazine/electronic versions of the mentioned sources) and summarize it in few sentences explaining why you chose that nutrition article and what information you like to emphasize or explain from that article
Answer #1

Colorectal cancer (CRC) is the fourth leading cause of death throughout the world. It is the third most commonly diagnosed cancer in the world, which accounts for about 1.4 million new cases diagnosed in 2012. The burden of colorectal cancer is expected to rise by 60% by 2030.

By the end of 2012, the number of people diagnosed with colorectal cancer in Australia was 52, 630 and it is the third most commonly diagnosed cancer in 2013 and is estimated to be the second most commonly diagnosed cancer by 2017 in Australia. There were about 14,962 new cases (8,214 males and 6,748 females) diagnosed in Australia in 2013. The number of new cases that are estimated to be diagnosed in Australia in 2017 is 16, 682. The average incidence rate reported in 2013 was 68 cases per 100,000 for males and 49 cases per 100,000 for females. The incidence was expected to increase with age for both genders.

A meta-analysis study indicating that red meat consumption is associated with cancer risk:

Summary:

Larsson et al conducted a meta-analysis of prospective studies on the consumption of red meat, processed meat, and colorectal cancer risk. In this, the authors studied 15 prospective studies on the consumption of red meat and 14 on processed meat. As per the study report, the relative risk values of colorectal cancer occurrence for the high to low meat intake categories were 1.20 (95% confidence interval (CI) = 1.15-1.42) and 28 (95% CI = 1.15-1.42) for processed meat and red meat respectively.

The estimated relative risks were 1.09 (95% CI = 1.05-1.13) for an increased consumption of processed meat more than 30 g/day and 1.28 (95% CI = 1.18-1.39) for an increased consumption of red meat more than 120 g/day. A positive association between the consumption of red meat, processed meat, and risk of colorectal cancer was observed in this study. A strong association between red meat consumption and incidence of rectal cancer was reported. Authors also estimated that reducing the maximum daily consumption of red meat to 65 g – 100 g per day can reduce the incidence of colorectal cancer.

Reference:

Larsson SC, Wolk A. Meat consumption and risk of colorectal cancer: a meta‐analysis of prospective studies. International journal of cancer. 2006 Dec 1;119(11):2657-64.